2a. Solve the quadratic equation

3z2 + (a - i)z + 3i = 0.


b. Solve the following system of equations of complex numbers:


z + iw = 1 + 2i

z - w = 1 -2i




1
Expert's answer
2022-02-23T12:51:23-0500

a.


"3z^2 + (a - i)z + 3i = 0."

"D=(a-i)^2-4(3)(3i)=a^2-2ai-1-36i"

"=a^2-1-2(a+18)i"

"z_1=\dfrac{-a+i-\sqrt{a^2-1-2(a+18)i}}{6}"

"z_2=\dfrac{-1+i+\sqrt{a^2-1-2(a+18)i}}{6}"

b.


"z + iw = 1 + 2i""z - w = 1 -2i"


"w(i+1) =4i""z(1+i)= 3+3i"

"w=\dfrac{4i}{1+i}""z=3"

"w=2+2i""z=3"